# Straggler-Resilient Decentralized Learning via Adaptive Asynchronous Updates

### v1

@@ Abstract
With the increasing demand for large-scale training of machine learning models, fully decentralized optimization methods have recently been advocated as alternatives to the popular parameter server framework. In this paradigm, each worker maintains a local estimate of the optimal parameter vector, and iteratively updates it by waiting and averaging all estimates obtained from its neighbors, and then corrects it on the basis of its local dataset. However, the synchronization phase is sensitive to stragglers. An efficient way to mitigate this effect is to consider asynchronous updates, where each worker computes stochastic gradients and communicates with other workers at its own pace. Unfortunately, fully asynchronous updates suffer from staleness of the stragglers’ parameters. To address these limitations, we propose a fully decentralized algorithm DSGD-AAU with adaptive asynchronous updates via adaptively determining the number of neighbor workers for each worker to communicate with. We show that DSGD-AAU achieves a linear speedup for convergence (i.e., convergence performance increases linearly with respect to the number of workers). Experimental results on a suite of datasets and deep neural network models are provided to verify our theoretical results.
00footnotetext: $*{}^{*}$qual contribution.
1 Introduction
@@
Highly over-parameterized deep neural networks (DNNs) have shown impressive results in a variety of machine learning (ML) tasks such as computer vision [he2016deep], natural language processing [vaswani2017attention] and many others. Its success depends on the availability of large amount of training data, which often leads to a dramatic increase in the size, complexity, and computational power of the training systems. In response to these challenges, the need for efficient parallel and distributed algorithms (i.e., data-parallel mini-batch stochastic gradient descent (SGD)) becomes even more urgent for solving large scale optimization and ML problems [bekkerman2011scaling, boyd2011distributed]. These ML algorithms in general use the parameter server (PS) [valiant1990bridging, smola2010architecture, dean2012large, ho2013more, li2014scaling, li2014communication] or ring All-Reduce [gibiansky17allreduce, goyal2017accurate] communication primitive to perform exact averaging on the local mini-batch gradients computed on different subsets of the data by each worker, for the later synchronized model update. However, the aggregation with PS or All-Reduce often leads to extremely high communication overhead, causing the bottleneck in efficient ML model training. Additionally, since all workers need to communicate with the server concurrently, the PS based framework suffers from the single point failure and scalability issues, especially in large-scale cloud computing systems [lian2017can].
@@
Recently, an alternative approach has been developed in ML community [lian2017can, lian2018asynchronous], where each worker keeps updating a local version of the parameter vector and broadcasts its updates only to its neighbors. This family of algorithms became originally popular in the control community, starting from the seminal work of [tsitsiklis1986distributed] on distributed gradient methods, in which, they are referred to as consensus-based distributed optimization methods or fully decentralized learning (DSGD) [^1] We interchangeably use “consensus-based methods” and “fully decentralized learning” (DSGD) in the rest of the paper. [neglia2020decentralized, neglia2019role]. The key idea of DSGD is to replace communication with the PS by peer-to-peer communication between individual workers.
@@
Most existing consensus-based algorithms assume synchronous updates, where all workers perform SGD at the same rate in each iteration. Such a synchronous update is known to be sensitive to stragglers, i.e., slow tasks, because of the synchronization barrier at each iteration among all workers. The straggler issues are especially pronounced in heterogeneous computing environments such as cloud computing, where the computing speed of workers often varies significantly. In practice, only a small fraction of workers participate in each training round, thus rendering the active worker (neighbor) set stochastic and time-varying across training rounds. An efficient way to mitigate the effect of stragglers is to perform fully asynchronous SGD, where each worker updates its local parameter immediately once it completes the local SGD and moves to the next iteration without waiting for any neighboring workers for synchronization [lian2018asynchronous, luo2020prague, assran2020asynchronous, wang2022asynchronous]. Unfortunately, such fully asynchronous SGD often suffers from staleness of the stragglers’ parameters and heavy communication overhead, and requires deadlock-avoidance (See Section 3). This raises the question: Is it possible to get the best of both worlds of synchronous SGD and fully asynchronous SGD in fully decentralized (consensus-based) learning?
@@
In this paper, we advocate adaptive asynchronous updates (AAU) in fully decentralized learning that is theoretically justified to keep the advantages of both synchronous SGD and fully asynchronous SGD. Specifically, in each iteration, only a subset of workers participate (rather than all as in synchronous SGD for decentralized learning), and each worker only waits for a subset of neighbors to perform local model updates (instead of waiting for no neighbor as in fully asynchronous SGD for decentralized learning). On the one hand, these participated workers are the fastest computational workers, which addresses the straggler issues in synchronous SGD. On the other hand, each worker does not need to perform local update immediately using stale information or communicate with all neighbors, which addresses the stateless and heavy-communication issues in fully asynchronous SGD. We make the following contributions:
@@
$\bullet$ We formulate the fully decentralized (consensus-based) optimization problem with AAU, and propose a decentralized SGD with AAU (DSGD-AAU) algorithm that adaptively determines the number of neighbor workers for each worker to perform local model update during the training time.
@@
$\bullet$ To our best knowledge, we present the first convergence analysis of fully decentralized learning with AAU that is cognizant of training process at each worker. We show that the convergence rate of DSGD-AAU on both independent-identically-distributed (i.i.d.) and non-i.i.d. datasets across workers is $\mathcal{O}\left(\frac{1}{\sqrt{NK}}\right)$, where $N$ is the total number of workers and $K$ is the total number of communication rounds. This indicates that DSGD-AAU achieves a linear speedup of communication w.r.t. $N$ for a sufficiently large $K.$ The state-of-the-art PSGD [lian2017can] and AD-PSGD [lian2018asynchronous] also achieve the same rate with synchronous updates or asynchronous updates with two randomly selected workers in each round, which lead to high communication costs and implementation complexity. In contrast we have the flexibility to adaptively determine the number of neighbor workers for each worker in the system. In addition, our proof does not require a bounded gradient assumption.
@@
$\bullet$ We conduct experiments on image classification (CIFAR-10, MNIST, Tiny-ImageNet) and next-character prediction (Shakespeare) tasks with several representative DNN models, demonstrating that DSGD-AAU substantially accelerates training of DNNs by mitigating the effect of stragglers and consistently outperforms state-of-the-art methods.
@@ 2 Background
$\min_{\mathbf{w}}~{}F({\mathbf{w}}):=\frac{1}{N}\sum\limits_{j=1}^{N}F_{j}({\mathbf{w}}),$ (1)
which aims to optimize a set of parameters ${\mathbf{w}}\in\mathbb{R}^{d\times 1}$ using $L$ examples from local training datasets ${\mathcal{D}}_{j}=\{{\mathbf{x}}_{\ell},\ell\in[L]\},\forall j\in[N]$. In particular, $F_{j}({\mathbf{w}})=\frac{1}{L}\sum_{{\mathbf{x}}_{\ell}\in{\mathcal{D}}_{j}}f({\mathbf{w}},{\mathbf{x}}_{\ell})$ with $f({\mathbf{w}},{\mathbf{x}}_{\ell})$ is the model error on the $\ell$-th element of dataset ${\mathcal{D}}_{j}$ when parameter ${\mathbf{w}}$ is used. In conventional decentralized learning, the distribution for each worker’s local dataset is usually assumed to be i.i.d.. Unfortunately, this assumption may not hold true in practice since data can be generated locally by workers based on their circumstances. Instead, we make no such assumption in our model and our analysis holds for both i.i.d. and non-i.i.d. local datasets across workers. See Section 4 for details.
@@
${\mathbf{w}}_{j}({k+1})=\sum_{i\in{\mathcal{N}}_{j}}\big{[}{\mathbf{w}}_{i}(k)-\eta g_{j}({\mathbf{w}}_{j}(k))\big{]}P_{i,j},$ (2)
where ${\mathbf{P}}=(P_{i,j})$ is a $N\times N$ non-negative matrix and we call it the consensus matrix, and $\eta>0$ is the learning rate. In other words, in each iteration, worker $j\in[N]$ computes a weighted average (i.e., consensus component) of the estimates of its neighbors and itself, and then corrects it by taking into account a stochastic subgradient $g_{j}({\mathbf{w}}_{j}(k))$ of its local function, i.e.,
$g_{j}({\mathbf{w}}_{j}(k)):=\frac{1}{|{\mathcal{C}}_{j}(k)|}\sum\limits_{{\mathbf{x}}_{\ell},\in{\mathcal{C}}_{j}(k)}\!\!\nabla f({\mathbf{w}}_{j}(k),{\mathbf{x}}_{\ell}),$ (3)
@@
(a) Decentralized SGD with synchronous updates.
@@
Figure 1: An illustrative example for decentralized SGD with (a) synchronous and (b) asynchronous updates in an arbitrary network topology with 4 workers. DSGD suffers from stragglers in (a), while workers in (b) suffer from staleness of the stragglers’ parameters, e.g., worker $4$ has updated its local parameter for $3$ times at iteration $k=7$, while worker $2$ has only updated once. Since worker $2$ may randomly pick worker $4$ for averaging as in AD-PSGD [lian2018asynchronous], worker $4$’s parameter update could be significantly dragged down by worker $2$’s stale information.
3 DSGD-AAU: Decentralized SGD with Adaptive Asynchronous Updates
The performance of synchronous updates in (2) may be significantly dragged down by stragglers (i.e., workers with slow computational capabilities), since every worker $j$ needs to wait for all neighbors $i\in{\mathcal{N}}_{j}$ to complete the current local gradient computation $g_{i}({\mathbf{w}}_{i}(k))$ at each iteration $k$ [lian2017can]. In other words, in each synchronous iteration, all workers participate, and each worker needs to wait for all its neighbors, as illustrated in Figure 1.
To mitigate the effect of stragglers in synchronous updates in (2), a possible way is to perform fully asynchronous updates, where each worker $j$ updates the local parameter ${\mathbf{w}}_{j}$ immediately once it completes the local gradient computation and moves to the next iteration [lian2018asynchronous, luo2020prague, assran2020asynchronous, wang2022asynchronous]. In other words, in each asynchronous iteration, only one (or more) worker participates, and each worker does not wait for any neighbors, as illustrated in Figure 1.
@@ 3 DSGD-AAU: Decentralized SGD with Adaptive Asynchronous Updates
For instance, the worker randomly selects a neighbor to exchange local parameters once it completes the local gradient computation [lian2018asynchronous], or the worker maintains a buffer to store information from neighbors whenever they complete a local update, and conducts a consensus update with the stale information in the buffer [assran2020asynchronous]. As a result, both approaches suffer from staleness of the stragglers’ parameters, as illustrated in Figure 1. Exacerbating the problem is the fact that the method in [lian2018asynchronous] only works for bipartite communication graph due to the deadlock issue, and the method in [assran2020asynchronous] suffers from heavy communications and memory space.
@@
To address aforementioned limitations, we advocate adaptive asynchronous updates (AAU) to get the best of both worlds of synchronous updates and fully asynchronous updates. Specifically, in each asynchronous iteration $k$, only a subset of ${\mathcal{N}}(k)\subset{\mathcal{N}}$ workers participate, and each worker $j$ only waits for a subset of neighbors, i.e., ${\mathcal{N}}_{j}(k)\subset{\mathcal{N}}_{j}$ to perform local model updates, where we have ${\mathcal{N}}(k)=\bigcup{\mathcal{N}}_{j}(k).$ We denote the cardinality of ${\mathcal{N}}(k)$ and ${\mathcal{N}}_{j}(k)$ as $a(k):=|{\mathcal{N}}(k)|<N$ and $a_{j}(k):=|{\mathcal{N}}_{j}(k)|<|{\mathcal{N}}_{j}|$, respectively. On the one hand, these $a(k)$ or $a_{j}(k)$ workers are the fastest computational workers, which addresses the straggler issues in synchronous updates. On the other hand, each worker does not need to perform local update immediately using stale information or communicate with all neighbors, which addresses the stateless and heavy-communication issues in fully asynchronous updates.
@@
As a result, each worker $j\in{\mathcal{N}}(k)$ first locally updates its parameter at iteration $k$ and then follows the consensus update only with workers from ${\mathcal{N}}_{j}(k)$, i.e.,
@@
${\mathbf{w}}_{j}(k)$ $=$ $\sum\nolimits_{i\in{\mathcal{N}}_{j}(k)}\tilde{{\mathbf{w}}}_{i}(k)P_{i,j}(k),$ (4)
@@
${\mathbf{W}}(k)$ $=[{\mathbf{W}}({k-1})-\boldsymbol{\eta}{\mathbf{G}}(k-1)]{\mathbf{P}}(k).$ (5)
The value of $a(k)$ and hence the consensus matrix ${\mathbf{P}}(k)$ in our AAU model is not static but dynamically changes over iterations by nature. We summarize the above procedure in Algorithm 1, and call the resulting decentralized SGD with adaptive asynchronous updates algorithm as DSGD-AAU.
Input: Topology ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, iterations $K$, and $\{{\mathbf{w}}_{j}(0)\}_{j=1}^{N}$.
@@
There are two open questions for DSGD-AAU. First, can DSGD-AAU still achieve a near-optimal convergence rate as in state-of-the-art decentralized method with synchronous updates [lian2017can] given that the analysis of asynchronous algorithm is already quite challenging? Second, how to determine $\mathcal{N}(k)$ (line 2 in Algorithm 1), and hence the number of fastest $a(k)$ workers at iteration $k$ to not only maintain the aforementioned best-of-both-world properties, but minimize the global training (convergence) time over the network? In the following, we provide an affirmative answer to the first question in Section 4, and then present a practical algorithm to determine $\mathcal{N}(k)$ (See Algorithm 2 for a fully decentralized implementation in Section 5).
4 Convergence Analysis
In this section, we analyze the convergence of DSGD-AAU. We show that a linear speedup for convergence (i.e., convergence performance increases linearly with respect to the number of workers) is achievable by DSGD-AAU. We first state our assumptions as follows.
4.1 Assumptions

### v2

@@ Abstract.
With the increasing demand for large-scale training of machine learning models, fully decentralized optimization methods have recently been advocated as alternatives to the popular parameter server framework. In this paradigm, each worker maintains a local estimate of the optimal parameter vector, and iteratively updates it by waiting and averaging all estimates obtained from its neighbors, and then corrects it on the basis of its local dataset. However, the synchronization phase is sensitive to stragglers. An efficient way to mitigate this effect is to consider asynchronous updates, where each worker computes stochastic gradients and communicates with other workers at its own pace. Unfortunately, fully asynchronous updates suffer from staleness of stragglers’ parameters. To address these limitations, we propose a fully decentralized algorithm DSGD-AAU with adaptive asynchronous updates via adaptively determining the number of neighbor workers for each worker to communicate with. We show that DSGD-AAU achieves a linear speedup for convergence and demonstrate its effectiveness via extensive experiments.
copyright: none
1. Introduction
@@
Deep neural networks (DNNs) have shown impressive results for a variety of machine learning (ML) tasks in the Internet of Things (IoT). Its success depends on the availability of large amount of training data, leading to a dramatic increase in the size, complexity and computational power of the IoT systems. In response to these challenges, the need for efficient parallel and distributed algorithms (i.e., data-parallel mini-batch stochastic gradient descent (SGD)) becomes urgent for solving large-scale optimization and ML problems (Bekkerman et al., 2011; Boyd et al., 2011). These ML algorithms in general use the parameter server (PS) (Smola and Narayanamurthy, 2010; Dean et al., 2012; Ho et al., 2013; Li et al., 2014) or ring All-Reduce (Gibiansky, 2017; Goyal et al., 2017) communication primitive to perform exact averaging on local mini-batch gradients computed on different subsets of the data by each worker, for the later synchronized model update. However, the aggregation with PS or All-Reduce often leads to extremely high communication overhead, causing the bottleneck for efficient ML model training in IoT. Additionally, since all workers need to communicate with the server concurrently, the PS based framework suffers from the single point failure and scalability issues, especially in IoT.
@@
Recently, an alternative approach has been developed in ML community (Lian et al., 2017, 2018), where each worker keeps updating a local version of the parameter vector and broadcasts its updates only to its neighbors. This family of algorithms became originally popular in the control community, starting from the seminal work of (Tsitsiklis et al., 1986) on distributed gradient methods, in which, they are referred to as consensus-based distributed optimization methods or fully decentralized learning (DSGD).The key idea of DSGD is to replace communication with the PS by peer-to-peer communication between individual workers.
@@
Most existing consensus-based algorithms assume synchronous updates, where all workers perform SGD at the same rate in each iteration. Such a synchronous update is known to be sensitive to stragglers, i.e., slow tasks, because of the synchronization barrier at each iteration among all workers. The straggler issues are especially pronounced in heterogeneous computing environments such as IoT, where the computing speed of workers often varies significantly. In practice, only a small fraction of workers participate in each training round, thus rendering the active worker (neighbor) set stochastic and time-varying across training rounds. An efficient way to mitigate the effect of stragglers is to perform fully asynchronous SGD, where each worker updates its local parameter immediately once it completes the local SGD and moves to the next iteration without waiting for any neighboring workers for synchronization (Lian et al., 2018; Luo et al., 2020; Assran and Rabbat, 2020; Wang et al., 2022). Unfortunately, such fully asynchronous SGD often suffers from staleness of the stragglers’ parameters and heavy communication overhead, and requires deadlock-avoidance (See Section 3). This raises the question: Is it possible to get the best of both worlds of synchronous SGD and fully asynchronous SGD in fully decentralized (consensus-based) learning?
@@
In this paper, we advocate adaptive asynchronous updates (AAU) in fully decentralized learning that is theoretically justified to keep the advantages of both synchronous SGD and fully asynchronous SGD. Specifically, in each iteration, only a subset of workers participate, and each worker only waits for a subset of neighbors to perform local model updates. On one hand, these participated workers are the fastest computational workers, which addresses the straggler issues in synchronous SGD. On the other hand, each worker does not need to perform local update immediately using stale information or communicate with all neighbors, which addresses the stateless and heavy-communication issues in fully asynchronous SGD. We make the following contributions:
@@
$\bullet$ We formulate the fully decentralized optimization problem with AAU, and propose a decentralized SGD with AAU (DSGD-AAU) algorithm that adaptively determines the number of neighbor workers to perform local model update during the training time.
@@
$\bullet$ To our best knowledge, we present the first convergence analysis of fully decentralized learning with AAU that is cognizant of training process at each worker. We show that the convergence rate of DSGD-AAU on both independent-identically-distributed (i.i.d.) and non-i.i.d. datasets across workers is $\mathcal{O}\left(\frac{1}{\sqrt{NK}}\right)$, where $N$ is the total number of workers and $K$ is the total number of communication rounds. This indicates that DSGD-AAU achieves a linear speedup of communication w.r.t. $N$ for a sufficiently large $K.$ This is the first linear speedup result for DSGD-AAU, and is highly desirable since it implies that one can efficiently leverage the massive parallelism in large-scale decentralized systems. The state-of-the-art PSGD (Lian et al., 2017) and AD-PSGD (Lian et al., 2018) also achieve the same rate with synchronous updates or asynchronous updates with two randomly selected workers in each round, which lead to high communication costs and implementation complexity. In contrast we have the flexibility to adaptively determine the number of neighbor workers for each worker in the system.
@@
$\bullet$ We conduct experiments on image classification and next-character prediction tasks with several representative DNN models, demonstrating that DSGD-AAU substantially accelerates training of DNNs by mitigating the effect of stragglers and consistently outperforms state-of-the-art methods.
2. Background
@@
(1) $\min_{\mathbf{w}}~{}F({\mathbf{w}}):=\frac{1}{N}\sum\limits_{j=1}^{N}F_{j}({\mathbf{w}}),$
which aims to optimize a set of parameters ${\mathbf{w}}\in\mathbb{R}^{d\times 1}$ using $L$ examples from local training datasets ${\mathcal{D}}_{j}=\{{\mathbf{x}}_{\ell},\ell\in[L]\},\forall j\in[N]$. In particular, $F_{j}({\mathbf{w}})=\frac{1}{L}\sum_{{\mathbf{x}}_{\ell}\in{\mathcal{D}}_{j}}f({\mathbf{w}},{\mathbf{x}}_{\ell})$ with $f({\mathbf{w}},{\mathbf{x}}_{\ell})$ is the model error on the $\ell$-th element of ${\mathcal{D}}_{j}$ when parameter ${\mathbf{w}}$ is used. In conventional decentralized learning, the distribution for each worker’s local dataset is assumed to be i.i.d.. Unfortunately, this assumption may not hold true in practice since data can be generated locally by workers based on their circumstances. Instead, we make no such assumption in our model and our analysis holds for both i.i.d. and non-i.i.d. local datasets across workers. See Section 4 for details.
@@
(2) ${\mathbf{w}}_{j}({k+1})=\sum_{i\in{\mathcal{N}}_{j}}\big{[}{\mathbf{w}}_{i}(k)-\eta g_{j}({\mathbf{w}}_{j}(k))\big{]}P_{i,j},$
where ${\mathbf{P}}=(P_{i,j})$ is a $N\times N$ non-negative matrix and we call it the consensus matrix, and $\eta>0$ is the learning rate. In other words, in each iteration, $j\in[N]$ computes a weighted average (i.e., consensus component) of the estimates of its neighbors and itself, and then corrects it by taking into account a stochastic subgradient $g_{j}({\mathbf{w}}_{j}(k))$ of its local function, i.e.,
(3) $g_{j}({\mathbf{w}}_{j}(k)):=\frac{1}{|{\mathcal{C}}_{j}(k)|}\sum\limits_{{\mathbf{x}}_{\ell},\in{\mathcal{C}}_{j}(k)}\!\!\nabla f({\mathbf{w}}_{j}(k),{\mathbf{x}}_{\ell}),$
@@
3. DSGD-AAU
The performance of synchronous updates in (2) may be significantly dragged down by stragglers since every worker $j$ needs to wait for all neighbors $i\in{\mathcal{N}}_{j}$ to complete the current local gradient computation $g_{i}({\mathbf{w}}_{i}(k))$ at each iteration $k$ (Lian et al., 2017). In other words, in each synchronous iteration, all workers participate, and each worker needs to wait for all its neighbors, as illustrated in Figure 1.
To mitigate the effect of stragglers in synchronous updates in (2), a possible way is to perform fully asynchronous updates, where each worker $j$ updates the local parameter ${\mathbf{w}}_{j}$ immediately once it completes the local gradient computation and moves to the next iteration (Lian et al., 2018; Luo et al., 2020; Assran and Rabbat, 2020; Wang et al., 2022). In other words, in each asynchronous iteration, only one (or more) worker participates, and each worker does not wait for any neighbors, as illustrated in Figure 1.
@@ 3. DSGD-AAU
Figure 1. An illustrative example for decentralized SGD with (a) synchronous and (b) asynchronous updates in an arbitrary network topology with 4 workers. DSGD suffers from stragglers in (a), while workers in (b) suffer from staleness of stragglers, e.g., worker $4$ has updated its local parameter for $3$ times at iteration $k=7$, while worker $2$ has only updated once. Since worker $2$ may randomly pick worker $4$ for averaging as in AD-PSGD (Lian et al., 2018), worker $4$’s parameter update could be significantly dragged down by worker $2$’s stale information.
@@
For instance, the worker randomly selects a neighbor to exchange local parameters once it completes the local gradient computation (Lian et al., 2018), or the worker maintains a buffer to store information from neighbors whenever they complete a local update, and conducts a consensus update with the stale information in the buffer (Assran and Rabbat, 2020). As a result, both approaches suffer from staleness of the stragglers’ parameters, as illustrated in Figure 1. Exacerbating the problem is the fact that the method in (Lian et al., 2018) only works for bipartite communication graph due to the deadlock issue, and the method in (Assran and Rabbat, 2020) suffers from heavy communications and memory space.
@@
To address the aforementioned limitations, we advocate adaptive asynchronous updates (AAU) to get the best of both worlds of synchronous updates and fully asynchronous updates. Specifically, in each asynchronous iteration $k$, only a subset of ${\mathcal{N}}(k)\subset{\mathcal{N}}$ workers participate, and each worker $j$ only waits for a subset of neighbors, i.e., ${\mathcal{N}}_{j}(k)\subset{\mathcal{N}}_{j}$ to perform local model updates, where we have ${\mathcal{N}}(k)=\bigcup{\mathcal{N}}_{j}(k).$ We denote the cardinality of ${\mathcal{N}}(k)$ and ${\mathcal{N}}_{j}(k)$ as $a(k):=|{\mathcal{N}}(k)|<N$ and $a_{j}(k):=|{\mathcal{N}}_{j}(k)|<|{\mathcal{N}}_{j}|$, respectively. On one hand, these $a(k)$ or $a_{j}(k)$ workers are the fastest computational workers, which addresses the straggler issues in synchronous updates. On the other hand, each worker does not need to perform local update immediately using stale information or communicate with all neighbors, which addresses the stateless and heavy-communication issues in fully asynchronous updates.
@@
As a result, $j\in{\mathcal{N}}(k)$ locally updates its parameter at iteration $k$ and then follows the consensus update only with workers in ${\mathcal{N}}_{j}(k)$,
@@
(4) ${\mathbf{w}}_{j}(k)$ $=$ $\sum\nolimits_{i\in{\mathcal{N}}_{j}(k)}\tilde{{\mathbf{w}}}_{i}(k)P_{i,j}(k),$
@@
(5) ${\mathbf{W}}(k)$ $=[{\mathbf{W}}({k-1})-\boldsymbol{\eta}{\mathbf{G}}(k-1)]{\mathbf{P}}(k).$
The value of $a(k)$ and the consensus matrix ${\mathbf{P}}(k)$ in our AAU model is not static but dynamically changes over iterations. We summarize the above procedure in Algorithm 1, and call the decentralized SGD with adaptive asynchronous updates algorithm as DSGD-AAU.
Input: Topology ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, iteration numbers $K$, and initialized $\{{\mathbf{w}}_{j}(0)\}_{j=1}^{N}$.
@@
There are two open questions for DSGD-AAU. First, can DSGD-AAU still achieve a near-optimal convergence rate as state-of-the-art decentralized method with synchronous updates (Lian et al., 2017) given that the analysis of asynchronous algorithm is already quite challenging? Second, how to determine $\mathcal{N}(k)$ (line 2 in Algorithm 1), and hence the number of fastest $a(k)$ workers at iteration $k$ to not only maintain the aforementioned best-of-both-world properties, but minimize the global training (convergence) time over the network? Note that at each iteration $k$ in Algorithm 1, the key is to determine the set of fastest workers, i.e., $\mathcal{N}(k)\in\mathcal{N}$, which have completed the local gradient computations, and only workers in $\mathcal{N}(k)$ actively conduct gossip-average with neighbors while workers in $\mathcal{N}\setminus\mathcal{N}(k)$ are still computing their local gradients. The iteration $k$ begins at finding $\mathcal{N}(k)$ and ends when $\mathcal{N}(k)$ is determined with workers inside it having completed the local gradient computations and the gossip-average operations. In the following, we provide an affirmative answer to the first question in Section 4, and then present a practical algorithm to determine $\mathcal{N}(k)$ (See Algorithm 2 for a fully decentralized implementation in Section 5).
4. Convergence Analysis
In this section, we analyze the convergence of DSGD-AAU.
4.1. Assumptions